i need to put this problem into 2 equal parts
5 7/8 into two equal pieces

Answer by jim_thompson5910(35256)

5 7/8 = 5+7/8

5 7/8 = 5(8/8)+7/8

5 7/8 = 40/8+7/8

5 7/8 = 47/8

-----------------------------

Now cut 47/8 in half by multiplying it by 1/2

47/8 * 1/2 = (47*1)/8*2) = 47/16

Then convert back to a mixed number

47/16 = 2 remainder 15

47/16 = 2 + 15/16

47/16 = 2 15/16

So 5 7/8 cut in half is 2 15/16